Which of the following is a consistent estimator:

(1) σ? - ΕΣ(Br-

(2) (BT - Â)

Where B1 ~ N(u,0?) . An estimator is consistent if it converges to the right answer as the sample size grows.
Answer #1

The solution to this problem is given below

Given Br ~ N(usa). The two estimators of or are given by h 3 (A) 22 +Š (OT-AY Now E 36-A)) [(:29] Now since 2) E ) - T- → E (

- By Cheby shers Inequality PERŽ%)* -8°/><]< von 63.0.) - Taking lim both sides we get, Timp 3 (8-48-09/2 c] - Lingerie =0 3

time [1(en A) - 0* | 2e] = 0 +(B-A) is a consistent a consistent is estimator. are constistent estimators .: Both estimators.